# Incorporating Coupling Knowledge into Echo State Networks for Learning Spatiotemporally Chaotic Dynamics

### v1

@@ II.3 Parallelized echo state network
Fig.1(a) illustrates an ordinary differential equation (ODE) system, which we will discuss further in Sec.III, while Fig. 1(b) depicts the corresponding Paral-ESN designed for learning its dynamics. We highlight the $i^{\mathrm{th}}$ reservoir computer and the corresponding dimensions of its input and predicted state with size $G=1$ in green, while the interacting dimensions of the input state with size $I=2$ are colored in blue. Note that the input state may consist of predicted states generated by the ESNs during the automatic prediction phase, where the dashed arrow identifies the feedback connection. This design pattern remains consistent across all reservoirs and is omitted here for brevity.
However, in Paral-ESN, local interactions are incorporated symmetrically only through the input layer. Moreover, ODE systems exhibit more substantial coupling structures, which can also serve as valuable guidance for model design. Building on this idea, we propose an ESN model that leverages prior physical knowledge of the coupling structure in an ODE system or the local interactions in a PDE system to guide the design of both the input and reservoir layers.
@@ III.1 Physical guidance in Dynamical Systems
where $D$ represents the ordinary derivative operator, $k$ is the order of the system, the index set $C_{i}=\{i,c_{i_{1}},\dots,c_{i_{l}}\}$ is a subset of $\{1,\dots,d_{u}\}$, and $l$ equals the number of its coupled dimensions besides itself, which could be 00. Fig. 1(a) schematically illustrates an observed ODE system with a coupled structure, which corresponds to the structure of the Lorenz-96 system—a benchmark DS used in this work. As the figure shows, the specific $i^{\mathrm{th}}$ dimension highlighted in green is coupled with its two preceding and one succeeding dimensions, which are colored in blue. Therefore, the index set for the coupled dimensions of the $i^{\mathrm{th}}$ dimension is given by $C_{i}=\{i-2,i-1,i,i+1\}$. The coupling for all other dimensions of this example system is identical and omitted here for brevity.
@@ III.2 Proposed model
In addition to the observed time series, suppose we also have prior knowledge of the coupling structure of the underlying DS. We can leverage this information as guidance to design a specialized ESN model, which we call the physics-guided clustered ESN (PGC-ESN). To illustrate the proposed approach, we assume that the underlying DS is governed by the ODE system shown in Fig. 1(a) and schematically illustrate the corresponding PGC-ESN in Fig. 1 (c). Specifically, our method involves constructing parallel, fully connected clusters within a single large reservoir. Each cluster corresponds to a particular dimension of the observed system and receives both its own dimension and the coupled dimensions of the input state. Additionally, each cluster is fully connected to other clusters, whose corresponding dimensions of the observed system are coupled with its own. The $i^{\mathrm{th}}$ dimension of the input and predicted states, along with the corresponding cluster in the reservoir, is highlighted in green, while the coupled dimensions and clusters are shown in blue. Unlike the Paralleled ESN method explored in Ref. Pathak et al., 2018a, which involves training each ESN separately, our proposed model maintains a single unified reservoir, enabling cohesive learning through a standard fully connected readout layer. This design methodology is consistently extended to PDE-based target systems, as illustrated in Appendix A.3.
@@
$\bm{r}_{i}(t+1)=f(\sum_{j\in C_{i}}\bm{W}^{\mathrm{in}}_{(i,j)}u_{j}(t)+\sum_{j\in C_{i}}\bm{W}_{(i,j)}\bm{r}_{j}(t)),$ (7)
where $\bm{r}_{i}(t)$ represents the state vector of the $i^{\mathrm{th}}$ cluster in the reservoir, and $u_{j}(t)$ denotes the $j^{\mathrm{th}}$ dimension of the input state at time $t$. The matrix $\bm{W}^{\mathrm{in}}_{(i,j)}$ contains the weight entries for connections from the $j^{\mathrm{th}}$ dimension of the input state to the $i^{\mathrm{th}}$ cluster in the reservoir, and $\bm{W}_{(i,j)}$ contains the weight entries for connections within the $i^{\mathrm{th}}$ cluster (if $j=i$) or from the $j^{\mathrm{th}}$ cluster to the $i^{\mathrm{th}}$ cluster in the reservoir (if $j\neq i$), where $j$ belongs to the set of coupled dimensions $C_{i}$ for the $i^{\mathrm{th}}$ dimension of the observed system. In closed-loop mode, the update equation remains almost the same, except that the reservoir is driven by the predicted state $\hat{\bm{u}}(t)$. A schematic illustration of the updation is provided in Appendix. A
@@ III.3 Comparison model without physical guidance
To evaluate whether the clustered structure in the reservoir layer can improve learning performance without incorporating the coupling structure of the observed system, we designed a model called the randomly clustered ESN (RandC-ESN). As shown in Fig. 1 (d), this model consists of a single reservoir composed of parallel, fully connected clusters. Each cluster corresponds to a specific dimension of the observed system and receives both its own dimension and a selection of randomly chosen dimensions from the input state, and the neuron connections between different clusters are randomly activated. We use a designed probability governing the random connections in both the input and reservoir layers to ensure that the overall sparsity of these layers matches that of the PGC-ESN.
@@ IV.1.1 Lorenz-96 system
$\frac{dx_{i}}{dt}=\alpha(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,$ (8)
for $i=1,\dots,N$, where $N\geq 4$, and $\alpha\geq 0$. The beginning and end are wrapped around such that $x_{-1}=x_{N-1}$, $x_{0}=x_{N}$ and $x_{N+1}=x_{1}$. This system consists of a nonlinear advective term $\alpha(x_{i+1}-x_{i-2})x_{i-1}$, a linear damping term $-x_{i}$, and an external forcing term $F$. The parameter $\alpha$ is added to control the nonlinearity of the system. When $\alpha=1$, the system is reduced to the standard Lorenz-96 model. In this work, we set $N=40$ and $F=8$. To simulate an observed trajectory, we start with a random initial state $\bm{u}_{0}=(x_{0},y_{0},z_{0})^{T}\sim\mathcal{N}(\mathbf{F},0.01\cdot\mathbf{I})$, where $\mathbf{I}$ is the identity matrix, and integrate it using the 4th-order Runge-Kutta method Butcher (2016) with step size $\Delta t=0.01$. We generate a trajectory of length $T=2\cdot 10^{5}$ by discarding the first $10^{5}$ samples to eliminate the effects of initial transients. The MLE for this system was calculated to be $\lambda\approx 1.69$ using the Shimoda-Nagashima method,Shimada and Nagashima (1979) and the corresponding Lyapunov time $T^{\lambda}\approx 0.59$.
@@ IV.1.2 Kuramoto–Sivashinsky equation
$\frac{\partial u}{\partial t}=-\frac{\partial^{4}u}{\partial x^{4}}-\left[1+\mu\cos\left(\frac{2\pi x}{\lambda}\right)\right]\frac{\partial^{2}u}{\partial x^{2}}-u\frac{\partial u}{\partial x},$ (9)
where $u$ is defined in the domain $x\in[0,L]$ with periodic boundary conditions, that is, $u(0,t)=u(L,t)$. The equation is spatially inhomogeneous when $\mu\neq 0$ and homogeneous when $\mu=0$, with both cases being considered in our experiments. The boundary size $L$ controls the dimensionality of the system’s attractor, scaling it linearly with a larger value. The system’s domain is spatially discretized to $d_{u}$ nodes with a fixed grid size, yielding an observable state $\bm{u}\in\mathbb{R}^{d_{u}}$. We solve the equation using the Exponential Time-Differencing Fourth-Order Runge–Kutta algorithm (ETDRK4) for stiff PDEs,Kassam and Trefethen (2005) with step size $\Delta t=0.25$. The simulated trajectory has a length of $T=2\cdot 10^{5}$, with the first $4\cdot 10^{4}$ samples being discarded. In our experiment, we set $L=100$, $d_{u}=256$ for both cases, and $\mu=0.1$, $\lambda=25$ for the inhomogeneous case. The calculated MLE values are $\lambda\approx 0.09$ for the homogeneous case and $\lambda\approx 0.08$ for the inhomogeneous case, with corresponding Lyapunov times of $T^{\lambda}\approx 11.02$ and $T^{\lambda}\approx 12.57$, respectively.
@@ IV.2 Evaluation metrics
In this work, we evaluate both geometric and temporal reconstruction using two metrics, referred to as the geometrical distance $D_{\mathrm{geom}}$ and the temporal distance $D_{\mathrm{temp}}$, Hess et al. (2023) for long-term attractor reconstruction, where smaller values indicate better reconstruction. Our aim was to capture the invariant geometrical structure of the system’s attractor in the state space and its temporal signature called the power spectrum. We provide the definitions of the two measures below, with further details on their numerical approximation in Appendix. B. Additionally, we assess the learning system’s short-term predictive performance using a metric called valid prediction time (VPT).Vlachas et al. (2020)
@@ IV.2.1 Geometrical distance
Geometrical distance Hess et al. (2023) is an invariant statistic used to measure the degree of overlap between true and reconstructed attractor geometries. Given the observed and predicted sequences of length $T$ on the invariant attractor, $\{\bm{u}(1),\dots\bm{u}(T)\}$ and $\{\hat{\bm{u}}(1),\dots\hat{\bm{u}}(T)\}$, which are generated by the ground truth system and the reconstructed system (i.e., the trained ML model), we can estimate the corresponding ideal invariant distributions $p(\bm{u})$ and $p(\bm{u}|\bm{\theta})$ over the state space, where $\bm{\theta}$ denotes the learned model parameters. The geometrical distance, $D_{\mathrm{geom}}$, is then defined as
$D_{\mathrm{geom}}\coloneqq D_{\mathrm{KL}}(p(\bm{u})\|p(\bm{u}|\bm{\theta}))=\int_{x\in\mathbb{R}^{N}}p(\bm{u})\log\frac{p(\bm{u})}{p(\bm{u}|\bm{\theta})}\,dx,$ (10)
where $D_{\mathrm{KL}}$ represents the Kullback–Leibler (KL) divergence between the two distributions.
@@ IV.2.2 Temporal distance
[…]
where $d_{u}$ is the dimension of the observed system’s state space, and $f_{i}$ and $g_{i}$ denote the power spectra for the $i^{\mathrm{th}}$ dimensions of the observed and predicted sequences, respectively.
@@ IV.2.3 Valid prediction time
To quantify the short-term predictive accuracy of the trained model over time, we need to compute the normalized root mean square error (NRMSE) Vlachas et al. (2020) which is given by
$\text{NRMSE}(\hat{\bm{u}}(t))=\sqrt{\left\langle\frac{(\hat{\bm{u}}(t)-\bm{u}(t))^{2}}{\bm{\sigma}^{2}(\bm{u}_{1:T})}\right\rangle},$ (12)
[…]
Another metric used to evaluate predictive accuracy is called the valid prediction time (VPT), Vlachas et al. (2020) which is the maximum time $t_{f}$ for which the model’s prediction maintains an NRMSE below a given threshold $\epsilon$, using the Lyapunov time $T^{\lambda}$ as the time unit. Formally, the VPT is given by
$\text{VPT}=\arg\max_{t_{f}}\{t_{f}\mid\text{NRMSE}(\hat{\bm{u}}(t))<\epsilon,\,\forall t\leq t_{f}\}.$ (13)
In this work, we set the threshold to $1$ and refer to it as VPT-1.0. This metric is particularly useful for assessing the model’s short-term predictive performance.
Figure 2: Predictive result for the Lorenz-96 system using a reservoir size of 500 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first two-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target sequence, the predicted sequence, and NRSE evolution across all dimensions, along with the NRMSE evolution averaged over all dimensions.

### v2

@@ II.3 Parallelized echo state network
Fig.1(a) illustrates an ordinary differential equation (ODE) system, which we will discuss further in Sec. III.1, while Fig. 1(b) depicts the corresponding Paral-ESN designed for learning its dynamics. We highlight the $i^{\mathrm{th}}$ reservoir computer and the corresponding dimensions of its input and predicted state with size $G=1$ in green, while the interacting dimensions of the input state with size $I=2$ are colored in blue. Note that the input state may consist of predicted states generated by the ESNs during the automatic prediction phase, where the dashed arrow identifies the feedback connection. This design pattern remains consistent across all reservoirs and is omitted here for brevity.
However, in Paral-ESN, local interactions are incorporated symmetrically only through the input layer. Moreover, ODE systems exhibit more substantial coupling structures, which can also serve as valuable guidance for model design. Building on this idea, we propose an ESN model that leverages physical knowledge of the coupling structure in an ODE system or the local interactions in a PDE system to guide the design of both the input and reservoir layers.
@@ III.1 Physical guidance in Dynamical Systems
where $D$ represents the ordinary derivative operator, $k$ is the order of the system, the index set $C_{i}=\{i,c_{i_{1}},\dots,c_{i_{l}}\}$ is a subset of $\{1,\dots,d_{u}\}$, and $l$ equals the number of its coupled dimensions besides itself, which could be 0. Fig. 1(a) schematically illustrates an observed ODE system with a coupled structure, which corresponds to the structure of the Lorenz-96 system—a benchmark DS used in this work. As the figure shows, the specific $i^{\mathrm{th}}$ dimension highlighted in green is coupled with its two preceding and one succeeding dimensions, which are colored in blue. Therefore, the index set for the coupled dimensions of the $i^{\mathrm{th}}$ dimension is given by $C_{i}=\{i-2,i-1,i,i+1\}$. The coupling for all other dimensions of this example system is identical and omitted here for brevity.
@@ III.2 Proposed model
In addition to the observed time series, suppose we also have prior knowledge of the coupling structure of the underlying DS. We can leverage this information as guidance to design a specialized ESN model, which we call the physics-guided clustered ESN (PGC-ESN). To illustrate the proposed approach, we assume that the underlying DS is governed by the ODE system shown in Fig. 1(a) and schematically illustrate the corresponding PGC-ESN in Fig. 1 (c). Specifically, our method involves constructing parallel, fully connected clusters within a single large reservoir. Each cluster corresponds to a particular dimension of the observed system and receives both its own dimension and the coupled dimensions of the input state. Additionally, each cluster is fully connected to the specific clusters whose corresponding dimensions of the observed system are coupled with its own. The $i^{\mathrm{th}}$ dimension of the input and predicted states, along with the corresponding cluster in the reservoir, are highlighted in green, while the coupled dimensions and clusters are shown in blue. Unlike the Paralleled ESN method explored in Ref. Pathak et al., 2018a, which involves training each ESN separately, our proposed model maintains a single unified reservoir, enabling cohesive learning through a standard fully connected readout layer. This design methodology is consistently extended to PDE-based target systems, as illustrated in Appendix A.2.
@@
$\bm{r}_{i}(t)=f(\sum_{j\in C_{i}}\bm{W}^{\mathrm{in}}_{(i,j)}u_{j}(t)+\sum_{j\in C_{i}}\bm{W}_{(i,j)}\bm{r}_{j}(t-1)),$ (7)
where $\bm{r}_{i}(t)$ represents the state vector of the $i^{\mathrm{th}}$ cluster in the reservoir, and $u_{j}(t)$ denotes the $j^{\mathrm{th}}$ dimension of the input state at time $t$. The matrix $\bm{W}^{\mathrm{in}}_{(i,j)}$ contains the weight entries for connections from the $j^{\mathrm{th}}$ dimension of the input state to the $i^{\mathrm{th}}$ cluster in the reservoir, and $\bm{W}_{(i,j)}$ contains the weight entries for connections within the $i^{\mathrm{th}}$ cluster (if $i=j$) or from the $j^{\mathrm{th}}$ cluster to the $i^{\mathrm{th}}$ cluster in the reservoir (if $i\neq j$), where $j$ belongs to the set of coupled dimensions $C_{i}$ for the $i^{\mathrm{th}}$ dimension of the observed system. In closed-loop mode, the update equation remains almost the same, except that the reservoir is driven by the predicted state $\hat{\bm{u}}(t)$. A schematic illustration of the updation is provided in Appendix A.3.
In practical scenarios, it is often necessary to consider settings where only time series data are available, and no prior knowledge of the underlying system’s coupling structure is provided. To address this, we propose a variant of PGC-ESN, referred to as AutoPGC-ESN, which is automatically guided by the extracted coupling information from the observed time series. Specifically, we estimate a coupling strength matrix $\bm{C}$, where each element $\bm{C}_{i,j}\geq 0$ quantifies the inferred coupling strength from the $j^{\mathrm{th}}$ dimension (sender) to the $i^{\mathrm{th}}$ dimension (receiver) of the underlying ODE system for $i\neq j$, and is defined as 0 for $i=j$. In our experiment, this estimation is performed by applying transfer entropy Schreiber (2000) to the pairwise sequences corresponding to dimensions $i$ and $j$ of the observed time series. Details regarding the definition and computation of transfer entropy are provided in Appendix B. For each dimension $i$, we define its inferred coupled dimensions set $\hat{C}_{i}$ as the union of $i$ and the $q$ dimensions with the highest (non-self) inferred coupling strength in the $i^{\mathrm{th}}$ row of the coupling strength matrix $\bm{C}$. The parameter $q$ is shared across all dimensions and affects the density of weight matrices in both the input and reservoir layers, as both densities now equal the ratio of $(1+q)$ to the dimension number of the underlying system. Using $\hat{C}_{i}$, we construct the coupling structure for the $i^{\mathrm{th}}$ cluster of AutoPGC-ESN’s reservoir in the same manner as in PGC-ESN. Notably, as the inferred coupling varies across dimensions, the resulting coupling structure of AutoPGC-ESN may differ across clusters accordingly. For PDE systems, a similar approach can be employed, with coupling defined and inferred at the level of spatial units as described in Sec. III.1.
@@ III.3 Comparison model without physical guidance
As an ablation study, we evaluate whether introducing the clustered structure in the reservoir layer—serving as the sole inductive bias—can improve learning performance without incorporating any spatial structure knowledge of the underlying system. To this end, we design a comparison model termed randomly clustered ESN (RandC-ESN). As shown in Fig. 1 (d), this model consists of a single reservoir composed of parallel, fully connected clusters. Each cluster corresponds to a specific dimension (or unit) of the observed ODE (or PDE) system and receives both its own dimension (unit) and a selection of randomly chosen dimensions (units) from the input state. Inter-cluster neuron connections are randomly activated without any weight bias, ensuring that no spatial structure knowledge of the underlying system is exploited. We use a designed probability governing the random connections in both the input and reservoir layers to ensure that the overall density of these layers matches that of the PGC-ESN.
@@ IV.1.1 Lorenz-96 system
$\frac{d\bm{u}_{i}}{dt}=\alpha(\bm{u}_{i+1}-\bm{u}_{i-2})\bm{u}_{i-1}-\bm{u}_{i}+F,$ (8)
for $i=1,\dots,d_{u}$, where the dimension $d_{u}\geq 4$, and $\alpha\geq 0$. The beginning and end are wrapped around such that $\bm{u}_{-1}=\bm{u}_{d_{u}-1}$, $\bm{u}_{0}=\bm{u}_{d_{u}}$ and $\bm{u}_{d_{u}+1}=\bm{u}_{1}$. This system consists of a nonlinear advective term $\alpha(\bm{u}_{i+1}-\bm{u}_{i-2})\bm{u}_{i-1}$, a linear damping term $\bm{u}_{i}$, and an external forcing term $F$. The parameter $\alpha$ is added to control the nonlinearity of the system. When $\alpha=1$, the system is reduced to the standard Lorenz-96 system. In this work, we set $d_{u}=40$ and $F=8$. To simulate an observed trajectory, we start with a random initial state $\bm{u}(0)\sim\mathcal{N}(\bm{F},\,(0.01)^{2}\mathbf{I})$, where $\bm{F}\in\mathbb{R}^{d_{u}}$ is a vector with all elements equal to $F$, and $\mathbf{I}\in\mathbb{R}^{d_{u}\times d_{u}}$ is the identity matrix. The system is then integrated using the 4th-order Runge-Kutta method Butcher (2016) with step size $\Delta t=0.01$. We generate a trajectory of length $T=2\cdot 10^{5}$ by discarding the first $10^{5}$ samples to eliminate the effects of initial transients. The MLE for this system was calculated to be $\lambda\approx 1.69$ using the Shimoda-Nagashima method, Shimada and Nagashima (1979) and the corresponding Lyapunov time $T^{\lambda}\approx 0.59$.
@@ IV.1.2 Kuramoto–Sivashinsky equation
$\frac{\partial u}{\partial t}=-\frac{\partial^{4}u}{\partial x^{4}}-\left[1+\mu\cos\left(\frac{2\pi x}{\gamma}\right)\right]\frac{\partial^{2}u}{\partial x^{2}}-u\frac{\partial u}{\partial x},$ (9)
where $u$ is defined on the domain $x\in[0,L]$ with periodic boundary conditions red, that is, $u(0,t)=u(L,t)$. The equation is spatially inhomogeneous when $\mu\neq 0$ and homogeneous when $\mu=0$, with both cases being considered in our experiments. The boundary size $L$ controls the dimensionality of the system’s attractor, scaling it linearly when $L$ is large. The system’s domain is spatially discretized to $d_{u}$ nodes with a fixed grid size, yielding an observable state $\bm{u}\in\mathbb{R}^{d_{u}}$. We solve the equation using the Exponential Time-Differencing Fourth-Order Runge–Kutta algorithm (ETDRK4) for stiff PDEs, Kassam and Trefethen (2005) with step size $\Delta t=0.25$. The simulated trajectory has a length of $T=2\cdot 10^{5}$, with the first $4\cdot 10^{4}$ samples being discarded to avoid initial transients. In our experiment, we set $L=100$, $d_{u}=256$ for both cases, and $\mu=0.1$, $\gamma=25$ for the inhomogeneous case. The calculated MLE values are $\lambda\approx 0.09$ for the homogeneous case and $\lambda\approx 0.08$ for the inhomogeneous case, with corresponding Lyapunov times of $T^{\lambda}\approx 11.02$ and $T^{\lambda}\approx 12.57$, respectively.
@@ IV.2 Evaluation metrics
In this work, we evaluate both geometric and temporal reconstruction using two metrics, referred to as the geometrical distance $D_{\mathrm{geom}}$ and the temporal distance $D_{\mathrm{temp}}$Hess et al. (2023) for long-term attractor reconstruction, where smaller values indicate better reconstruction. Our aim was to capture the invariant geometrical structure of the system’s attractor in the state space and its temporal signature called the power spectrum. We provide the definitions of the two measures below, with further details on their numerical approximation in Appendix C. Additionally, we assess the learning system’s short-term predictive performance using a metric called valid prediction time (VPT). Vlachas et al. (2020)
@@ IV.2.1 Geometrical distance
Geometrical distance Hess et al. (2023) is an invariant statistic used to measure the degree of overlap between true and reconstructed attractor geometries. Given the observed and predicted sequences of length $T_{\mathrm{pred}}$ on the invariant attractor, $\{\bm{u}(1),\dots\bm{u}(T_{\mathrm{pred}})\}$ and $\{\hat{\bm{u}}(1),\dots\hat{\bm{u}}(T_{\mathrm{pred}})\}$, which are generated by the ground truth system and the reconstructed system (i.e., the trained ML model), we can estimate the corresponding ideal invariant distributions $p(\bm{u})$ and $p(\bm{u}|\bm{\theta})$ over the state space, where $\bm{\theta}$ denotes the learned model parameters. The geometrical distance $D_{\mathrm{geom}}$ is then defined as
$D_{\mathrm{geom}}\coloneqq D_{\mathrm{KL}}(p(\bm{u})\|p(\bm{u}|\bm{\theta}))=\int_{\bm{u}\in\mathbb{R}^{d_{u}}}p(\bm{u})\log\frac{p(\bm{u})}{p(\bm{u}|\bm{\theta})}\,d\bm{u},$ (10)
where $D_{\mathrm{KL}}$ represents the Kullback–Leibler (KL) divergence between the two distributions. Fig. 2(a) shows both the target and predicted attractors projected onto the first three-dimensional subspace in the evaluation of ESN models on the Lorenz-96 system. The $D_{\mathrm{geom}}$ values reported in Table 2 quantify the discrepancy between the target and predicted attractors for each model, averaged across sequences and model initializations.
@@ IV.2.2 Temporal distance
$D_{\mathrm{temp}}\coloneqq\frac{1}{d_{u}}\sum_{i=1}^{d_{u}}H(f_{i},g_{i})=\frac{1}{d_{u}}\sum_{i=1}^{d_{u}}\sqrt{1-\int_{-\infty}^{\infty}\sqrt{f_{i}(\omega)g_{i}(\omega)}\,d\omega},$ (11)
where $d_{u}$ is the dimension of the observed system’s state space, and $f_{i}$ and $g_{i}$ denote the power spectra over frequency $\omega$ for the $i^{\mathrm{th}}$ dimensions of the observed and predicted sequences, respectively. Fig. 2(b) shows both the target and predicted power spectra in the evaluation of ESN models on the Lorenz-96 system. The $D_{\mathrm{temp}}$ values reported in Table 2 quantify the discrepancy between the power spectra of the target and predicted sequences for each model, averaged across sequences and model initializations.
@@ IV.2.3 Valid prediction time
To quantify the short-term predictive accuracy of the trained model over time, we need to compute the normalized root mean square error (NRMSE) Vlachas et al. (2020), which is given by
$\text{NRMSE}(\hat{\bm{u}}(t))=\sqrt{\left\langle\frac{(\hat{\bm{u}}(t)-\bm{u}(t))^{2}}{\bm{\sigma}^{2}(\bm{u}_{1:T_{\mathrm{ts}}})}\right\rangle},$ (12)
where $\bm{u}(t)$ and $\hat{\bm{u}}(t)$ are the observed and predicted states at time-step $t$, $\bm{u}_{1:T_{\mathrm{ts}}}$ is the observed training sequence of length $T_{\mathrm{ts}}$, and $\bm{\sigma}(\bm{u}_{1:T_{\mathrm{ts}}})$ denotes the vector of standard deviations over time, computed per state-space dimension of this sequence. The notation $\langle\cdot\rangle$ represents the average over all state-space dimensions. Without this averaging, we obtain the normalized root square error (NRSE), which is a vector spanning all state-space dimensions. These two values will be used to plot the evolution of the model’s predictive accuracy.
Another metric used to evaluate predictive accuracy is called the valid prediction time (VPT), Vlachas et al. (2020) which is the maximum time $t_{m}$ for which the model’s prediction maintains an NRMSE below a given threshold $\epsilon$, using the Lyapunov time $T^{\lambda}$ as the timescale. Formally, the VPT is given by
$\text{VPT}=\max_{t_{m}}\{t_{m}\mid\text{NRMSE}(\hat{\bm{u}}(t))<\epsilon,\,\forall t\leq t_{m}\}.$ (13)
In this work, we set the threshold to $\epsilon=1$ and refer to this metric as VPT-1. This metric is particularly useful for assessing the model’s short-term predictive performance.
Figure 2: Predictive result for the Lorenz-96 system using a reservoir size of 500 for PGC-ESN, Paral-ESN, RandC-ESN, and Std-ESN (from left to right). (a) The attractor on the first three-dimensional subspace. (b) The power spectrum. (c) Contour plots of the target and predicted sequences (d) NRMSE evolution and contour plots of the NRSE evolution.